OpenStudy (anonymous):

i saw a proof in wikipedia about 0.9999999........... the proof says that the above is equal to 1 . now if f(x) = [x] , where [ ] represents greatest integer function would the limit exist at x=1

OpenStudy (anonymous):

in fact .9 repeating IS 1

OpenStudy (anonymous):

do u want a proof?

OpenStudy (anonymous):

oh that is not what you are asking i see

OpenStudy (anonymous):

and the answer is no the limit at 1 of the greatest integer function does not exist

OpenStudy (anonymous):

yes..certainly..the limit at an integer is nonexistent

OpenStudy (anonymous):

the greatest integer function has a jump at 1.

OpenStudy (anonymous):

then if x = 0.99999999999........ would d limit exist then

OpenStudy (anonymous):

no it doesnt exist

OpenStudy (anonymous):

because 0.99999999999999999999999 IS infact 1

OpenStudy (anonymous):

i see the problem. you are assuming that .999... is somehow less than one. it is if you truncate, but if you write .9 repeating that number is 1 and no other

OpenStudy (anonymous):

yes satellites right..0.9 bar is infact one

OpenStudy (anonymous):

and no other

OpenStudy (anonymous):

if technically seeing it approaches to 1

OpenStudy (anonymous):

which in fact is a very good explanation of why the limit does not exist! [.9] = [.99] = [.999]=[.9999]=....=0

OpenStudy (anonymous):

no matter how far out you go, as long as you stop then you will get 0 for the greatest integer function. but evaluated at .9 repeating = 1 you get 1

OpenStudy (anonymous):

the real question is what does it mean to say .9 repeating =1, and it really comes down to "what does .9 repeating actually mean?"

OpenStudy (anonymous):

yes sir ....

OpenStudy (anonymous):

here is an example you are probably more familiar with. most people probably have seen that \[\frac{1}{3}=.333...\] yes?

OpenStudy (anonymous):

see if u convert 1/9 into a decimal...u get 0.111111111111111111 so 1/9= 0.111111111111...... multiply both sides by 9 1/9 x 9 = 0.1111111111111111111111111111111111....... x 9 1= 0.999999....

OpenStudy (anonymous):

@him this is an explanation, but it begs the question of what .1 repeating actually means. after all if i cannot parse .1 repeating i cannot multiply it by 9, since i do not have enough time

OpenStudy (anonymous):

.1 repeating actually means 1/9..u can see that in the method to convert a repeating decimal into a rational number representation...

OpenStudy (anonymous):

point being that i do not have enough time to divide 1 by 9! i will be doing it forever. i see the pattern but that does not explain the meaning of .1 repeating or .3 repeating or .9 repeating

OpenStudy (anonymous):

x = 0.111... 10x = 1.111... subtract one from two... 9x = 1 x = 1/9 hence 0.111...=1/9

OpenStudy (anonymous):

i know the proof sir . i actually want to ask will the limit exist at x = 0.9999999999..... when f(x) = [x]

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

sir mat bula yaar..itna buddha bhi nahin hoon

OpenStudy (anonymous):

i can make no sense out of writing an infinite number of decimals without resorting to a limit statement. \[\bar{.1}=lim_{n\rightarrow \infty}\sum_{k=1}^n\frac{1}{10^k}\]

OpenStudy (anonymous):

\[\bar{.9}=lim_{n\rightarrow \infty}\sum_{k=1}^n\frac{9}{10^k}\]

OpenStudy (anonymous):

its a GP

OpenStudy (anonymous):

i cannot add subtract multiply or divide an infinite number of decimals. i have to say precisely what i mean by them. this requires work, although the "trick" of multiplying and subtracting is what we will do. but the understanding means i have to kn ow what a limit is and how i can say .9 repeat in IS 1

OpenStudy (anonymous):

both my proof and urs are on wikipedia....it says its been proven by methods of varying mathematical rigour....

OpenStudy (anonymous):

sure of course. you have to begin where you are

OpenStudy (anonymous):

if i want to convince someone that .9 repeating is 1, that is exactly what i would say, what you said!

OpenStudy (anonymous):

but that begs the question of what .9 repeating MEANS. that is my point.

OpenStudy (anonymous):

it means one... "In mathematics, the repeating decimal 0.999..., which may also be written as 0.9, or 0.(9), denotes a real number that can be shown to be the number one. In other words, the symbols 0.999... and 1 represent the same number. "

OpenStudy (anonymous):

this says nothing other than .9 repeating is 1. it does not explain what it means write in infinite repeating decimal

OpenStudy (anonymous):

its just a fraction which when evaluated gives a nonterminating repeating sequence of digits..dont get technical now,.. : )

OpenStudy (anonymous):

there is nothing special about decimal fractions. you can use base anything you like. so for example i can use base 2 and then .1 repeating = 1 or i can use base 3 and then .2 repeating = 1

OpenStudy (anonymous):

or use base 4 and get .4 repeating =1 you get the picture

OpenStudy (anonymous):

that was wrong. in base 4 it is .3 repeating = 1

OpenStudy (anonymous):

so in base 4 .1 repeating would mean \[\bar{.1}=lim_{n\rightarrow \infty}\sum_{k=1}^n\frac{1}{4^k}\]

OpenStudy (anonymous):

true....

OpenStudy (anonymous):

so now wht do u want to prove?

OpenStudy (anonymous):

only point i am trying to make, a simple point really, is that we need to understand what we mean by a repeating decimal. it is a limit of a sum. we are so used to writing numbers in base ten that we forget that they have a specific meaning.

OpenStudy (anonymous):

right....

OpenStudy (anonymous):

even terninating decimals right? i have to remind myself that \[.251=\frac{2}{10}+\frac{5}{100}+\frac{1}{1000}\]

OpenStudy (anonymous):

so for greatest integer function we see that \[ f(\bar{.9})=f(lim_{n\rightarrow \infty}\sum_{k=1}^n\frac{9}{10^k})=1\]

OpenStudy (anonymous):

is f(x) 1 or is the limit 1?

OpenStudy (anonymous):

but \[f(\sum_{k=1}^n\frac{9}{10^k})=0\] for all finite n

OpenStudy (anonymous):

i think the limit is 1...so for all mathematical puposes 0.999.. is 1

OpenStudy (anonymous):

yes i wrote limit. take the limit, get 1. evaluate greatest integer function, get 1. but until you pass to the limit you do not get 1 and therefore the greatest integer function will be 0

OpenStudy (anonymous):

so just before it the limit is 0, bt just after it it is 1..so it doesnt exist

OpenStudy (anonymous):

and thus not continuous of course. picture tells us that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well there is no "just before the limit". there is the limit and there is everything else. but yes, until you take the limit you get something less than one. and once you take the limit you get 1

OpenStudy (anonymous):

by just before i mean the left hand limit..r u phd in maths?

OpenStudy (anonymous):

btw if the greatest integer function were continuous you would be able to "take the limit outside the function" as they say. which of course you cannot do

OpenStudy (anonymous):

actually no i am just a well trained dog. that is the beauty of the internet http://www.unc.edu/depts/jomc/academics/dri/idog.jpg

OpenStudy (anonymous):

nice...lol